OpenStudy (anonymous):

Adam can wash and detail his car in 2 hours, and Lisa can wash and detail her car in 4 hours. If they work together, how long will it take them to wash and detail both cars?

OpenStudy (jamesj):

The way to think about problems like this is in terms of \( \it rates \). The rate at which Michael can wash his car is \[ R_M = 1/2 \ car/hour \] In 1 hour, he washes 1/2 a car. Lisa's rate is \[ R_L = 1/4 \] So far so good?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

so together they could do 3/4 of a car in 1 hour?

OpenStudy (jamesj):

Hence if they work together their combined rate is \[ R = R_M + R_L \] \[ = \frac{1}{2} + \frac{1}{4} = \frac{3}{4} \] Therefore how long does it take them together to wash one car?

OpenStudy (anonymous):

so would it be 2 2/3 hours to wash both cars? lol i kinda think i skipped to the end

OpenStudy (jamesj):

To wash one car it takes \[ t = \frac{1 \ car}{ cars/hour} = \frac{1}{3/4} = \frac{4}{3} \ hours \] Hence two cars is 8/3 hours, which is what you wrote down. It's worth however understanding this mathematics, because if you understand this once through and through, these problems will become trivial in the future.

OpenStudy (anonymous):

ahh ok i think i got it now..